OpenStudy (anonymous):

Assume a 150 pound person burns 120 calories after 2 hours of sleep and 240 calories after 4 hours of sleep. Part 1 (2 points): Write an equation to represent this data. Use calories as the y-coordinate and hours as the x-coordinate. Part 2 (2 points): What is the slope of the equation found in part 1? What does the slope represent in terms of the information provided? Part 3 (2 points): How many calories will this person burn after 8 hours of sleep? Using complete sentences, explain how the equation, slope, or graph can help to predict calories burned.

OpenStudy (anonymous):

Part 1: Create a table to find the two point require to make a line. \[\begin{array}{|c|c|} \hline hours & calories \\ \hline x & y\\ \hline 2 & 120 \\\hline 4 & 240 \\ \hline \end{array}\] Slope is defined as the change in x over the change in y or: \(\Large{\frac{\triangle x}{\triangle y} = \frac{x_2-x_1}{y_2-y_1} = \frac{4-2}{240-120} = \frac{2}{120} = \frac{1}{60} =} \) Slope. Use the point-slope formula to find the equation of the line. \(y – y1 = m(x – x1)\) and plug in your values for x and y and (slope or m). \(y-120 = \frac{1}{60}(x-2)\) Solve for y if required by your teacher. Part 2: What does the slope mean? For every 1 hour of sleep, how many calories are burned? Part3: How many calories burned after 8 hours? Plug in the number and solve the equation.